The density of gold at 20°C is 19,300kg/m3. Find the density at 100°C?

1
Expert's answer
2021-11-21T17:25:13-0500

"V=V_0(1+\alpha\Delta T)\to\rho=\rho_0/(1+\alpha\Delta T)\to"


"\rho=19300/(1+42\cdot10^{-6}\cdot(100-20))=19235\ (kg/m^3)"